# Deep Graph Anomaly Detection: A Survey and New Perspectives

### v1

@@ 4.2.1 Feature Interpolation
GraphSMOTE [150], GraphMixup [126] and GraphENS [94] are representative methods that utilize feature interpolation to address the challenge of biased GAD (C4 in Sec. 2.2). These methods are designed for imbalanced node classification, which can be considered as supervised GAD in a closed-set setting since they do not consider the detection of novel/unknown anomaly types that are not illustrated by the labeled training anomaly examples [92, 90, 121, 16, 161, 1]. In particular, GraphSMOTE [150] extends SMOTE to synthesize new nodes and edges in graph data. Different from generic data, generating new nodes in a graph requires the connections of these nodes to existing nodes. Thus, an edge generator is trained simultaneously in GraphSMOTE to model the relations between existing and new nodes when using the SMOTE-style approach to generate the new nodes. A similar work is done in AugAN [159] that performs interpolation on the feature representations of the nodes that are similar to labeled anomalies to increase the number of training anomaly examples. However, directly performing the interpolation may produce out-of-domain samples due to the extreme sparsity of the minority classes. To alleviate this issue, GraphMixup [94] is introduced to construct semantic relation spaces that allow the interpolation to be performed at the semantic level. In GraphENS [94], the synthesized nodes are created using an adaptive interpolation rate that is determined by the distance between the minority and majority class nodes, and its neighborhood is built in a stochastic manner based on the distance between the ego-network nodes of a minority node and a target node. BAT [72] generates virtual nodes for each class as “shortcuts” connecting to the other nodes based on posterior likelihoods, where the feature representation of the generated nodes is generated based on a feature interpolation operation.
@@ 4.2.2 Noise Perturbation
Advantages. (i) The graph instance generation may allow the synthesis of novel anomalies that facilitate the identification of unseen anomalies. (ii) Noise perturbation generation can create more diversified training samples, making the trained models less sensitive to small changes in the input data.
Disadvantages. (i) The generated instances can inevitably include some out-of-domain examples, which may deviate from the optimization objective, rendering the detection models less effective. (ii) Determining the optimal level of noise in the noise perturbation can be challenging and may require extensive experimentation.
@@
5 Proxy Task Design
@@
5.1 Graph Reconstruction
Data reconstruction aims to learn low-dimensional feature representations of data for reconstructing given data instances, which is widely used in tabular data, and image/video data to detect anomalies [91]. Given that the normal samples often occupy most of the dataset, it helps guarantee that the representations can retain the normal information, and thus, normal samples will be reconstructed with a smaller reconstruction error than anomaly samples. As a commonly used technique in this category, Graph autoencoder (GAE) [41] consists of a graph encoder and a graph decoder and it is easy to implement. Formally, given a graph with $\bf{X}$ and ${\bf{A}}$, the graph reconstruction can be formulated as
@@ 5.1 Graph Reconstruction
where the GNN encoder and decoder are parameterized by ${\Theta_{Enc}}$ and ${\Theta_{Dec}}$ respectively, $\bf{Z}$ are the representations of the nodes in the latent space. ${\bf{\hat{X}}}$ and ${\widehat{\bf{A}}}$ are the reconstructed attributes and graph structure respectively. The optimization objective of the methods in this group can be unified as follows
@@
5.2 Graph Contrastive Learning
Graph contrastive learning (GCL) aims to learn effective representations without human annotated labels so that similar graph instances are pulled together in the representation space, while dissimilar instances are far apart. A principled GCL framework for GAD is to devise a self-supervised contrastive learning task with suitable positive and negative pairs to learn the underlying dominant (normal) patterns in graph-structured data. The framework generally first utilizes a GNN network to learn the representations of graph instances, and then it employs a contrastive loss function $L_{\text{c}}$ to maximize the similarity of the positive views while minimizing the similarity of the negative pairs [118]:
@@ 5.2 Graph Contrastive Learning
where $R$ is the number of node-subgraph pairs sampled during inference to obtain a stable anomaly score. This framework inspires a number of follow-up methods, including the use of prior knowledge of different anomaly types to generate the negative samples [135], supervised positive/negative subgraph pairs [157], and multi-view/scale subgraph generation [51, 155, 27, 132].
The above methods are focused on node-level anomaly detection in static graph data. Contrastive learning is also used in dynamic GAD or graph-level anomaly detection. For example, TADDY [70] applies a dynamic graph transformer that aggregates spatial and temporal knowledge simultaneously to learn the representations of edges. It specifies ${\bf{s}}_{i}$ and ${\tilde{\mathbf{s}}_{i}}$ by constructing the positive edge using the existing edges in the training set and generates the anomalous edges via negative sampling. SIGNET [67] is designed for graph-level anomaly detection, which first constructs two different views using dual hypergraph transformation and then maximizes the mutual information between the bottleneck subgraph from two views. The estimated mutual information can be used to evaluate the graph-level abnormality.
@@
5.3 Graph Representation Distillation
@@
This category of methods learns the representation from the teacher model initialized by a GNN. Subsequently, a GNN-based student network is trained to replicate the representation outputs of the teacher model. The GNN-based teacher and student networks are formulated as follows
@@
GlocaKD [79] is an early framework, in which the teacher model $GN{N_{{\rm{teacher}}}}$ is implemented using a random GCN. Then the distillation function $KD(\cdot)$ is instantiated using KL divergence to minimize the graph- and node-level prediction errors of the representations yielded by the random GCN. The anomaly score in GlocaKD is defined as the prediction error at the graph and node levels. To support better distillation of the normal graph representations, several distillation models have been proposed for GAD with new architectures. For example, the dual-student-teacher model, called GLADST, consists of one teacher model and two student models [59], in which the teacher model, trained with a heuristic loss, is designed to make the representations more divergent. Unlike the traditional teacher-student model, GLADST trains the two student models on normal and abnormal graphs separately to capture the normality and abnormality better. Unlike GlocalKD that uses a random GNN to be the teach network, the approach FGAD uses a pre-trained anomaly detector as the teacher model [9]. Then the student model is designed with a graph isomorphism network (GIN) backbone and a projection head to improve the robustness of GAD.
@@
Challenge Addressed. GNN-enabled knowledge distillation enhances the ability of addressing graph structure-aware GAD. (C1). By distilling knowledge from a well-trained teacher model, the student model gains improved generalization across different GAD scenarios, showcasing enhanced robustness in GAD (C5).
5.4 Adversarial Graph Learning
@@
It is worth mentioning that the purpose of generation is different from the noise perturbation generation in the generative GNNs in Sec. 4.2. The former mainly uses the graph structure information and interpolation operations in the latent space to generate new node representations without using adversarial learning process, while the latter is focused on generating node representations from a prior distribution through the adversarial learning to learn the latent normality.
@@
This group of methods typically employs GNNs to learn the representations of graph instances and a generator network for generating graph instances based on a prior. Formally, these methods follow the following framework:
@@
5.5 Score Prediction
This group of methods focuses on how to make full use of labeled data to build an end-to-end anomaly score prediction model. Unlike approaches that directly apply GNNs for classification with labeled abnormal and normal nodes, this method is designed for the scenarios where only some graph instances are known to be normal and abnormal instances.
Assumption. The anomaly scores of normal graph instances follow a prior distribution while that for abnormal instances significantly deviate from the distribution.
@@ 5.5 Score Prediction
6 Graph Anomaly Measures
This category of methods aims to discuss GAD methods that focus on designing anomaly measures for evaluating the abnormality of instances in the graph. These methods generally perform anomaly scoring by incorporating some key abnormal graph characteristics into deep graph learning methods. As shown in Fig. 6, they can be generally divided into four categories, including one-class distance measure, local affinity measure, community adherence measure, and graph isolation-based approaches.
@@ 6 Graph Anomaly Measures
6.1 One-class Classification Measure
The one-class classification measure refers to evaluating the distance between each instance and a one-class center of the instances for anomaly scoring [105]. This method can also be applied to graph data, where the GNN is typically trained to minimize the volume of a hypersphere that encloses the representations of the graph instances. The key intuition is that anomalous graph instances differ significantly from the normal ones, causing them to fall outside the hypersphere that encompasses most of the normal graph instances.
@@ 6.2 Community Adherence
This community adherence-based method first leverages a mapping function to learn the representation of graph instances. A clustering or community discovery method is then applied to group the graph instances. Since anomalies generally exhibit significantly weaker community adherence, the anomaly score can be defined by the minimum distance to the centers of the communities.
@@
This approach is analogous to clustering-based anomaly detection in non-graph data [89], but here it needs to capture the graph characteristics for GAD. To this end, MHGL [158] utilizes GNNs and a multi-hypersphere learning objective to learn multiple groups of fine-grained normal patterns, enclosing each group using a corresponding hypersphere in the latent space while simultaneously pushing labeled anomalies far away from these hyperspheres. The anomaly score is defined as the Euclidean distance between a test instance and the nearest hypersphere center. Netwalk [140] is a method for both anomalous node and edge detection where k-means clustering was applied to group the existing node/edge into different groups in a feature representation space. The anomaly score of node/edge is measured as its distance to the center of its closest cluster.
@@
6.3 Local Affinity
There are many graph properties that can be important for GAD, such as connectivity, degree distribution, and clustering coefficient. Local affinity is a graph property that integrates multiple properties for evaluating the normality and abnormality of graph instances. The affinity may be defined in various ways, such as the number of connections of a graph instances to neighboring instances and clustering coefficient of the connected subgraphs. The key intuition is that the normal graph instances typically have a strong affinity with its neighbors, whereas an anomalous instance has a significantly weaker affinity with its neighbors. Thus, the local affinity can serve as the inverse of anomaly score.
@@ 6.3 Local Affinity
TAM [98] is a seminal work that introduces local affinity as an anomaly measure. It aims to learn tailored node representations for GAD by maximizing the local affinity of nodes to their neighbors. It is optimized on truncated graphs where non-homophily edges are removed iteratively to mitigate its adverse effects on the local affinity measure. The learned representations result in a significantly stronger local affinity for the normal nodes than the abnormal nodes. CLAD [53] instead measures the affinity based on the discrepancy between a node and its neighbors using Jenson-Shannon Divergence. The anomaly score is obtained from the affinity for each node in terms of both graph structure and attributes. PREM [85] eliminates the message-passing propagation in the regular GNNs by using an ego neighbor matching-based contrastive learning module. It aims to learn discriminative patterns between the local ego network and the neighboring instances. These GAD methods are focused on the anomaly score of a node based on its affinity to its neighboring nodes. Exploring beyond the node-level affinity can be one potential approach for subgraph- or graph-level anomaly detection.
@@ 6.4 Graph Isolation
The methods in this group need to first learn the representations of the instances using a graph encoder $GNN_{enc}$:
@@
Challenge Addressed. The graph isolation measure can adapt traditional anomaly measures to GAD with the power of GNN-based representation learning (C1). The isolation mechanism is highly efficient, allowing good scalability to GAD on large graphs (C2).
7 Research Opportunities
@@
Advanced Graph Anomaly Measures. Most current GAD methods are built upon proxy tasks or focused on the GNN backbones using traditional non-graph anomaly measures, as summarized in Table II. Consequently, they may fail to learn feature representations that encapsulate holistic consideration of graph structure and attributes specifically for GAD. Therefore, it is crucial to devise anomaly measures that go beyond traditional anomaly measures and proxy tasks, such as local node affinity [98], for developing more dedicated methods for GAD.
@@
GAD on Complex Graphs. Most existing GAD methods are focused on small-scale (e.g., less than millions of nodes/edges), static, or homogeneous graph data, as shown in Tables III and IV in Appendix B. However, many real-world graphs can involve millions/billions of heterogeneous nodes/edges [56], such as real-life citation networks, social networks, and financial networks. The nodes/edges may appear in a streaming fashion, where the models can access to only limited graph data at one time step and may need to adapt to new normal/abnormal patterns as the graph evolves [103]. Current methods may be adapted to handle these graphs, but their performance would become less effective since their primary design do not consider those complexity. GAD methods designed for graphs with two or more of these complexities are required.
@@
Handling Anomaly Camouflage and Contamination. In GAD, anomalies might easily hide their abnormal characteristics by mimicking the structure and attributes of their neighboring instances. There have been some approaches for addressing this problem, e.g., via selecting relevant features, incorporating domain knowledge, or using adversarial training [25, 99], but they often rely on the prior knowledge about what specific features the attackers may use in the camouflage. A related important problem is the issue of anomaly contamination in the training data. Current methods are mostly unsupervised, working on anomaly-contaminated training data, but the anomalous instances in the graph can largely bias the message-passing in the GNNs [98], leading to less expressive feature representations. Recent approaches, such as semi-supervised GAD on a small set of labeled normal graph instances [99] or training GNNs using truncated graph data [98], may offer effective methodologies for handling these issues.
@@
Interpretable GAD. As shown by the summarized detection performance results in Tables V, VI and VII in Appendix C, current GAD methods have shown impressive success in detecting anomalous graph instances, but they generally ignore the interpretability of their detection results. In addition to accurate detection, interpretable GAD also requires an explanation about why a graph instance is identified as anomalous within a given graph structure, making it different from explaining anomalies in non-graph data. Exploring information such as local graph structure, human feedback, and/or domain knowledge [67] would be some interesting directions for providing the structure-aware anomaly explanation. Also, the obtained anomaly explanation may in turn be further leveraged to improve the detection performance.
@@
Open-set Supervised GAD. As shown in Table II, there have been many supervised GAD methods, most of which essentially tackle an imbalanced binary classification problem. Such formulation is often questionable since anomalies per se can draw from very different distributions and cannot be treated as from one concrete class. Open-set supervised GAD also trains the detectors with labeled normal and anomalous examples (i.e., seen anomalies), but it assumes an open set of anomaly classes (i.e., there are anomaly classes that are not illustrated by the training anomaly samples) rather than the closed-set assumption in most existing studies [121]. Thus, it is a more realistic supervised GAD setting. Methods for this setting have shown significantly better performance than unsupervised and fully supervised methods on visual data [1, 16, 136, 161] and tabular data [91, 92, 90]. Recent studies in this line also show similar advantages on graph data [121, 158]. Exploring better modeling of the normal patterns while fitting the seen anomalies could be an effective approach to avoid overfitting of the seen anomalies (i.e., reducing misclassification of the unseen anomalies as normal).
@@
Foundation Models for GAD. Leveraging foundation models for downstream tasks has been emerging as one effective direction to empower the sample-efficient performance in the downstream tasks, including graph-related tasks [130, 111, 62, 112], owing to their superior generalization ability. Two main directions for the GAD task involve the training of graph foundation models (GFMs) for GAD and the exploitation of large language models (LLMs) for GAD. There have been a number of successful tuning of foundation models for anomaly detection on image data [162, 156, 50] and video data [129, 128, 127, 107] via proper prompt crafting, prompt learning, or in-context learning. Similar approaches may be explored for GAD, such as the in-context learning-based GAD method in [68] inspiring from [162]. This new paradigm is plausible for GAD in several aspects, such as zero/few-shot detection on target graph data, inductive GAD (most existing GAD methods take the transductive approach), and interpretable GAD with text description.
8 Conclusion

### v2

@@ IV-B1 Feature Interpolation
GraphSMOTE [152], GraphMixup [128] and GraphENS [96] are representative methods that utilize feature interpolation to address the challenge of biased GAD (C4 in Sec. II-B). These methods are designed for imbalanced node classification, which can be considered as supervised GAD in a closed-set setting since they do not consider the detection of novel/unknown anomaly types that are not illustrated by the labeled training anomaly examples [94, 92, 123, 16, 163, 1]. In particular, GraphSMOTE [152] extends SMOTE to synthesize new nodes and edges in graph data. Different from generic data, generating new nodes in a graph requires the connections of these nodes to existing nodes. Thus, an edge generator is trained simultaneously in GraphSMOTE to model the relations between existing and new nodes when using the SMOTE-style approach to generate the new nodes. A similar work is done in AugAN [161] that performs interpolation on the feature representations of the nodes that are similar to labeled anomalies to increase the number of training anomaly examples. However, directly performing the interpolation may produce out-of-domain samples due to the extreme sparsity of the minority classes. To alleviate this issue, GraphMixup [96] is introduced to construct semantic relation spaces that allow the interpolation to be performed at the semantic level. In GraphENS [96], the synthesized nodes are created using an adaptive interpolation rate that is determined by the distance between the minority and majority class nodes, and its neighborhood is built in a stochastic manner based on the distance between the ego-network nodes of a minority node and a target node. BAT [73] generates virtual nodes for each class as “shortcuts” connecting to the other nodes based on posterior likelihoods, where the representation of the generated nodes is generated based on a feature interpolation operation.
@@ IV-B2 Noise Perturbation
Advantages. (i) The graph instance generation may allow the synthesis of novel anomalies that facilitate the identification of unseen anomalies. (ii) Noise perturbation generation can create more diversified training samples, making the trained models less sensitive to small changes in the input data. Disadvantages. (i) The generated instances can inevitably include some out-of-domain examples, which may deviate from the optimization objective, rendering the detection models less effective. (ii) Determining the optimal level of noise in the noise perturbation can be challenging and may require extensive experimentation.
@@
V Proxy Task Design
@@
V-A Graph Reconstruction
Data reconstruction aims to learn low-dimensional feature representations of data for reconstructing given data instances, which is widely used in tabular data, and image/video data to detect anomalies [93]. Given that the normal samples often occupy most of the dataset, it helps guarantee that the representations can retain the normal information, and thus, normal samples will be reconstructed with a smaller reconstruction error than anomaly samples. As a commonly used technique in this category, Graph autoencoder (GAE) [41] consists of a graph encoder and a graph decoder and it is easy to implement the reconstruction process. As shown in Fig. 5(a), given a graph with $\bf{X}$ and ${\bf{A}}$, the graph reconstruction can be formulated as
@@ V-A Graph Reconstruction
where the GNN encoder and decoder, which typically follow the general form of the FAG mechanism in Eq. (2), are designed to obtain the feature representation for each node by aggregating the representations of its neighbors. They are parameterized by ${\Theta_{Enc}}$ and ${\Theta_{Dec}}$ respectively. $\bf{Z}$ are the representations of the nodes in the latent space. ${\bf{\hat{X}}}$ and ${\widehat{\bf{A}}}$ are the reconstructed attributes and graph structure respectively. The optimization objective of the methods in this group can be unified as follows
@@
V-B Graph Contrastive Learning
Graph contrastive learning (GCL) aims to learn effective representations without human annotated labels so that similar graph instances are pulled together in the representation space, while dissimilar instances are far apart. A principled GCL framework for GAD is to devise a self-supervised contrastive learning task with suitable positive and negative pairs to learn the underlying dominant (normal) patterns in graph-structured data. As shown in Fig. 5(b), the framework generally first utilizes a GNN network to learn the representations of graph instances and build the positive pair and negative pair based on the structure, and then employs a contrastive loss function $L_{\text{c}}$ to maximize the similarity of the positive views while minimizing the similarity of the negative pairs [121]:
@@ V-B Graph Contrastive Learning
where $R$ is the number of node-subgraph pairs sampled during inference to obtain a stable anomaly score. This framework inspires a number of follow-up methods, including the use of prior knowledge of different anomaly types to generate the negative samples [137], supervised positive/negative subgraph pairs [159], and multi-view/scale subgraph generation [52, 156, 26, 134]. The above methods are focused on node-level anomaly detection in static graph data. Contrastive learning is also used in dynamic GAD or graph-level anomaly detection. For example, TADDY [71] applies a dynamic graph transformer that aggregates spatial and temporal knowledge simultaneously to learn the representations of edges. It specifies ${\bf{s}}_{i}$ and ${\tilde{\mathbf{s}}_{i}}$ by constructing the positive edge using the existing edges in the training set and generates the anomalous edges via negative sampling. SIGNET [68] is designed for graph-level anomaly detection, which first constructs two different views using dual hypergraph transformation and then maximizes the mutual information between the bottleneck subgraph from two views. The estimated mutual information can be used to evaluate the graph-level abnormality.
@@
V-C Graph Representation Distillation
@@
As shown in Fig. 5 (c), this category of methods learns the representation from a teacher model initialized by a GNN. Subsequently, a GNN-based student network is trained to replicate the representation outputs of the teacher model. The GNN-based teacher and student networks are formulated as follows
@@
GlocaKD [80] is an early framework, in which the teacher model $GN{N_{{\rm{teacher}}}}$ is implemented using a random GCN. Then the distillation function $KD(\cdot)$ is instantiated using KL divergence to minimize the graph- and node-level prediction errors of the representations yielded by the random GCN. The anomaly score in GlocaKD is defined as the prediction error at the graph and node levels. To support better distillation of the normal graph representations, several distillation models have been proposed for GAD with new architectures. For example, the dual-student-teacher model, called GLADST, consists of one teacher model and two student models [60], in which the teacher model, trained with a heuristic loss, is designed to make the representations more divergent. Unlike the traditional teacher-student model, GLADST trains the two student models on normal and abnormal graphs separately to capture the normality and abnormality better. Unlike GlocalKD that uses a random GNN to be the teach network, the approach FGAD uses a pre-trained anomaly detector as the teacher [9]. Then the student is designed with a graph isomorphism network (GIN) and a projection head to improve the robustness of GAD.
@@
Challenge Addressed. GNN-enabled knowledge distillation enhances the ability of addressing graph structure-aware GAD. (C1). By distilling knowledge from a well-trained teacher, the student gains improved generalization across different GAD scenarios, showcasing enhanced robustness in GAD (C5).
V-D Adversarial Graph Learning
@@
It is worth mentioning that the purpose of generation is different from the noise perturbation generation in the generative GNNs in Sec. IV-B. The former mainly uses the graph structure information and interpolation operations in the latent space to generate new node representations without using adversarial learning process, while the latter is focused on generating node representations from a prior distribution through the adversarial learning to learn the latent normality.
@@
This group of methods (see Fig. 5 (d)) typically employs GNNs to learn the representations of graph instances and a generator network for generating graph instances based on a prior. The discriminator is utilized to distinguish whether a graph instance comes from the generator or the original data. Formally, these methods follow the following framework:
@@
V-E Score Prediction
The score prediction based-methods focuses on how to make full use of labeled data to build an end-to-end anomaly score prediction model, as shown in Fig. 5 (e). Unlike approaches that directly apply GNNs for classification with labeled abnormal and normal nodes, this method is designed for the scenarios where only some graph instances are known to be normal and abnormal instances.
Assumption. The anomaly scores of normal graph instances follow a prior distribution while those for abnormal instances significantly deviate from the distribution.
@@ V-E Score Prediction
VI Graph Anomaly Measures
This category of methods aims to discuss GAD methods that focus on designing anomaly measures for evaluating the abnormality of instances in the graph. These methods generally perform anomaly scoring by incorporating some key abnormal graph characteristics. As shown in Fig. 6, they can be generally divided into four categories, including one-class distance measure, local affinity measure, community adherence measure, and graph isolation-based approaches.
@@ VI Graph Anomaly Measures
VI-A One-class Classification Measure
One-class classification measures refer to evaluating the distance between each instance and a one-class center of the instances for anomaly scoring [108]. This method can be applied to graph data, with the GNNs trained to minimize the volume of a hypersphere that encloses the representations of the graph instances. As illustrated in Fig. 6 (a), the key intuition is that anomalous graph instances differ significantly from the normal ones, causing them to fall outside the hypersphere that encompasses most of the normal graph instances.
@@ VI-B Community Adherence
This type of methods first leverages a mapping function to learn the representation of instances. A clustering or community discovery method is then applied to group the graph instances, as shown in Fig. 6 (b), where the instances are grouped into three distinct clusters. Since anomalies generally exhibit significantly weaker community adherence, the anomaly score can be defined by the minimum distance to the centers of the communities.
@@
This approach is analogous to clustering-based anomaly detection in non-graph data [91], but here it needs to capture the graph characteristics for GAD. To this end, MHGL [160] utilizes GNNs and a multi-hypersphere learning objective to learn multiple groups of fine-grained normal patterns, enclosing each group using a corresponding hypersphere in the latent space while simultaneously pushing labeled anomalies far away from these hyperspheres. The anomaly score is defined as the Euclidean distance between a test instance and the nearest hypersphere center. Netwalk [142] is a method for both anomalous node and edge detection where k-means clustering was applied to group the existing node/edge into different groups. The anomaly score of node/edge is measured as its distance to the center of its closest cluster.
@@
VI-C Local Affinity
There are many graph properties that can be important for GAD, such as connectivity, degree distribution, and clustering coefficient. Local affinity is a graph property that integrates multiple properties for evaluating the normality and abnormality of graph instances. As shown in Fig. 6 (c), the affinity may be defined in various ways, such as the number of connections of a graph instance to neighboring instances, the similarity with neighbors, or the clustering coefficient of the connected subgraphs. The key intuition is that the normal graph instances typically have a strong affinity with their neighbors, whereas an anomalous instance has a significantly weaker affinity with its neighbors. Thus, the local affinity can serve as the inverse of the anomaly score.
@@ VI-C Local Affinity
TAM [101] is a seminal work that introduces local affinity as an anomaly measure. It aims to learn tailored node representations for GAD by maximizing the local affinity of nodes to their neighbors. It is optimized on truncated graphs where non-homophily edges are removed iteratively to mitigate its adverse effects on the local affinity measure. The learned representations result in a significantly stronger local affinity for the normal nodes than the abnormal nodes. CLAD [54] instead measures the affinity based on the discrepancy between a node and its neighbors using Jenson-Shannon Divergence. The anomaly score is obtained from the affinity for each node in terms of both graph structure and attributes. PREM [87] eliminates the message-passing propagation in regular GNNs by using an ego neighbor matching-based contrastive learning module. It aims to learn discriminative patterns between the local ego network and the neighboring instances. These GAD methods are focused on the anomaly score of a node based on its affinity to its neighboring nodes. Similar ideas have also been explored in very recent prompt tuning or in-context learning methods for GAD [86, 69, 100]. Exploring beyond the node-level affinity can be one potential approach for subgraph- or graph-level anomaly detection.
@@ VI-D Graph Isolation
As shown in Fig. 6 (d), the methods in this group need to first learn the representations of the instances using a graph encoder $GNN_{enc}$:
@@
Challenge Addressed. This measure can adapt traditional anomaly measures to GAD with the power of GNN-based representation learning (C1). The isolation mechanism is highly efficient, allowing good scalability on large graphs (C2).
VII Research Opportunities
@@
Advanced Graph Anomaly Measures. Most GAD methods are built upon proxy tasks or focused on the GNN backbones using traditional non-graph anomaly measures, as summarized in Table 1 in Appendix A. Consequently, they may fail to learn feature representations that encapsulate holistic graph structure and attributes specifically for GAD. Therefore, it is crucial to devise anomaly measures that go beyond traditional anomaly measures and proxy tasks, such as local node affinity [101], for developing more dedicated methods for GAD.
@@
GAD on Complex Graphs. Most existing GAD methods are focused on small-scale (e.g., less than millions of nodes/edges), static, or homogeneous graph data, as shown in Tables 2 and 3 in Appendix B. However, many real-world graphs can involve millions/billions of heterogeneous nodes/edges [57], such as real-life citation networks, social networks, and financial networks. The nodes/edges may appear in a streaming fashion, where the models can access to only limited graph data at one time step and may need to adapt to new normal/abnormal patterns as the graph evolves [106]. Current methods may be adapted to handle these graphs, but their performance would become less effective since their primary design do not consider those complexity. GAD methods designed for graphs with two or more of these complexities are required.
@@
Handling Anomaly Camouflage and Contamination. In GAD, anomalies might easily hide their abnormal characteristics by mimicking the structure and attributes of their neighboring instances. There have been some approaches for addressing this problem, e.g., via selecting relevant features, incorporating domain knowledge, or using adversarial training [24, 102], but they often rely on the prior knowledge about what specific features the attackers may use in the camouflage. A related problem is anomaly contamination in the training data. Current methods are mostly unsupervised, working on anomaly-contaminated training data, but the anomalous instances in the graph can largely bias the message-passing [101], leading to less expressive representations. Recent approaches, such as semi-supervised GAD on a small set of labeled normal instances [102] or training GNNs using truncated graph data [101], may offer effective methodologies for handling these issues.
@@
Interpretable GAD. As shown by the summarized detection performance results in Tables 4, 5, and 6 in Appendix C, current GAD methods have shown impressive success in detecting anomalous graph instances, but they generally ignore the interpretability of their detection results. In addition to accurate detection, interpretable GAD also requires an explanation about why a graph instance is identified as anomalous within a given graph structure, making it different from explaining anomalies in non-graph data. Exploring information such as local graph structure, human feedback, and/or domain knowledge [68] would be some interesting directions for providing the structure-aware anomaly explanation. Also, the obtained anomaly explanation may in turn be further leveraged to improve the detection performance.
@@
Open-set Supervised GAD. As shown in Table 1 in Appendix A, there have been many supervised GAD methods, most of which essentially tackle an imbalanced binary classification problem. Such formulation is often questionable since anomalies per se can draw from very different distributions and cannot be treated as from one concrete class. Open-set supervised GAD also trains the detectors with labeled normal and anomalous examples (i.e., seen anomalies), but it assumes an open set of anomaly classes (i.e., there are anomaly classes that are not illustrated by the training anomaly samples) rather than the closed-set assumption in most existing studies [123]. Thus, it is a more realistic supervised GAD setting. Methods for this setting have shown significantly better performance than unsupervised and fully supervised methods on visual data [1, 16, 138, 163] and tabular data [93, 94, 92]. Recent studies also show similar advantages on graph data [123, 160]. Exploring better modeling of the normal patterns while fitting the seen anomalies could be an effective approach to avoid overfitting of the seen anomalies (i.e., reducing misclassification of the unseen anomalies as normal).
@@
Foundation Models for GAD. Leveraging foundation models (FMs) for downstream tasks has been emerging as one effective direction to empower the sample-efficient performance in the downstream tasks, including graph-related tasks [132, 114, 63, 115], owing to their superior generalization ability. Two main directions for the GAD task involve the training of graph foundation models (GFMs) for GAD and the exploitation of large language models (LLMs) for GAD. There have been a number of successful tuning of FMs for anomaly detection on image data [164, 157, 51] and video data [131, 130, 129, 110] via proper prompt crafting, prompt learning, or in-context learning. Similar approaches may be explored for GAD. For example, inspired by [164], an in-context residual learning-based method is explored for few-shot GAD [69]; general abnormality/normality graph patterns and prompts are learned for zero/few-shot cross-dataset GAD in [86, 100]. This new paradigm is plausible for GAD in several aspects, such as zero/few-shot detection on target graph data, inductive GAD, and interpretable GAD with text description.
VIII Conclusion